y for a police academy, candidates must score in the top 18% on a general es test. Assume test scores are normally distributed with a mean of 210 anda 3. To qualif abiliti standard deviation of 15. Find the lowest possible score to qualify. Draw the curve and shade the region. 5 210
Answer #1

Z-score of top 18% is \small Z=0.915

Given \small \mu =210 and \small \sigma =15

Let the qualifying lowest score be X

\small X=\mu +Z\sigma =210+0.915\times 15=223.725\approx 224

165 180 195 210 225 240 255